# Advances in Appfl: A Comprehensive and Extensible Federated Learning Framework

### v1

@@ Abstract
Federated learning (FL) is a distributed machine learning paradigm enabling collaborative model training while preserving data privacy. In today’s landscape, where most data is proprietary, confidential, and distributed, FL has become a promising approach to leverage such data effectively, particularly in sensitive domains such as medicine and the electric grid. Heterogeneity and security are the key challenges in FL, however; most existing FL frameworks either fail to address these challenges adequately or lack the flexibility to incorporate new solutions. To this end, we present the recent advances in developing Appfl, an extensible framework and benchmarking suite for federated learning, which offers comprehensive solutions for heterogeneity and security concerns, as well as user-friendly interfaces for integrating new algorithms or adapting to new applications. We demonstrate the capabilities of Appfl through extensive experiments evaluating various aspects of FL, including communication efficiency, privacy preservation, computational performance, and resource utilization. We further highlight the extensibility of Appfl through case studies in vertical, hierarchical, and decentralized FL. Appfl is open-sourced at https://github.com/APPFL/APPFL.
@@ I Introduction
Availability of extensive training data is becoming increasingly crucial for developing more capable machine learning (ML) models, especially as these models continue to grow in size and complexity. Nonetheless, most of the data in today’s landscape is confidential and distributed across various data silos [1]. This distribution makes it difficult to collect the data for centralized model training, posing significant challenges in fully leveraging the existing data to train more powerful ML models. In this context, federated learning (FL), a distributed ML paradigm, offers a promising solution to utilize data from multiple data owners without direct data sharing [2, 3, 4, 5].
@@
In FL, multiple data owners, referred to as clients, collaborate under a central server to train a shared ML model by iterating the following two steps: (1) each client trains an ML model using its local dataset and submits the updated model to the server, and (2) the server aggregates these local models to update the global model and then sends it back to the clients for further local training. In this way, FL leverages data from multiple data sources to build a more powerful and robust model without requiring data centralization, thereby protecting data privacy. Consequently, FL has been widely adopted in domains such as medicine [6, 7, 8], finance [9, 10], and electric grid [11], where data privacy is paramount. Depending on the amount, capability, and availability of client devices, FL is broadly categorized into two types, cross-device FL and cross-silo FL [5]. In cross-device FL, numerous mobile or IoT devices with limited computing power and intermittent availability collaboratively train relatively small models such as keyboard suggestion and hot word detection models [12, 13, 14]. In contrast, cross-silo FL involves fewer but more reliable and powerful clients, typically represented by large data silos and institutions, to develop more complex ML models with extensive parameters.
@@
While FL can be conceptually simplified to traditional machine learning with an additional global aggregation operation, its distributed nature introduces significant challenges in terms of heterogeneity and security when adopted for real-world applications. Data heterogeneity, stemming from the unbalanced, and non-independent and identically distributed (non-IID) nature of client local datasets, can lead to varied local training objectives across clients and potentially degrade the performance of the global model [15]. Additionally, the heterogeneity in computation and communication, caused by diverse computing capabilities and network connectivity of client devices, can severely impact the efficiency of FL training [16, 17]. This is particularly problematic in synchronous FL algorithms, where the server has to wait for all clients to submit their local models before global aggregation. With regard to security, FL is vulnerable to various attacks. Untrusted clients might maliciously attack FL experiments by submitting corrupted local models, and there is also a risk that training data could be reconstructed from the model updates sent by clients, thereby compromising data privacy [5, 18].
@@
Most existing FL frameworks, such as Flower [19], FedML [20], and FedScale [21], do not adequately address the full spectrum of challenges inherent in FL. For example, some frameworks do not support asynchronous aggregation that could improve training efficiency, lack implementations of robust authentication and privacy preservation algorithms, or fail to offer user-friendly interfaces for developers to easily integrate new solutions or algorithms. To bridge these gaps, we developed the Advanced Privacy-Preserving Federated Learning (Appfl) framework, a comprehensive and extensible FL framework that builds on and improves the work presented in [22]. Appfl supports both single-node and multi-node simulations, as well as distributed deployment of FL experiments. It features advanced aggregation strategies to address data heterogeneity [15] and various asynchronous aggregation strategies to boost training efficiency in environments with computation heterogeneity [23]. Additionally, Appfl incorporates versatile communication protocols [24], data transfer methods [25], and compression strategies to meet different communication requirements and enhance communication efficiency [26]. It also includes robust authentication via Globus [27, 28, 29], along with plugins for adding new authentication methods, and implements privacy preservation strategies to prevent the reconstruction of training data [30]. Moreover, Appfl is extensible; it follows a modular design that enables users and developers to seamlessly adapt the framework for different use cases and integrate custom algorithmic solutions to tackle various FL challenges.
@@
  • Advance Appfl, an open-source FL framework for both FL users and developers that provides established solutions to common FL challenges for FL users and offers flexible and modular interfaces that facilitate easy integration of new algorithmic solutions for FL developers
@@ II-A Heterogeneity in Federated Learning
Heterogeneity is one of the significant challenges in FL due to its distributed nature across multiple clients [5, 31, 32]. This heterogeneity can be categorized into three primary types: data heterogeneity, computation heterogeneity, and communication heterogeneity.
@@
Data heterogeneity arises from the fact that client datasets are often unbalanced and non-IID, meaning they may not be representative of the overall population distribution. This discrepancy leads to varying local training objectives among clients, causing their locally trained models to diverge from one another, a phenomenon known as client drift [33]. As a result, simple weighted averaging of local models, as in the FedAvg strategy [3], may degrade the performance of the global model as data heterogeneity increases [15]. Several solutions have been proposed to address this issue on both the server and client sides. For instance, server-side optimizations such as FedAvgM [15], FedAdam, FedAdagrad, and FedYogi [34] have been introduced to enhance FL performance on non-IID data. Client-side approaches such as FedProx [16] and SCAFFOLD [33] incorporate correction terms into the client’s local objective function to reduce excessive drift between local and global models.
@@
Computation heterogeneity occurs when the computing devices of FL clients have varying computing power, resulting in large variants in the local training times [35]. This variance poses challenges for synchronous FL strategies, where the server must wait for all clients to submit their local models before proceeding with global aggregation. Such delays from slower clients can reduce training efficiency and lead to underutilization of computing resources. In order to address this issue, various asynchronous aggregation strategies have been proposed. These strategies, including FedAsync [36], FedASO [37], FedBuff [38], AREA [39], and FedCompass [23], update the global model immediately upon receiving models from one or a few clients. These methods are particularly beneficial in environments with heterogeneous computing capabilities because they minimize client idle time and enhance resource utilization. Other approaches include disregarding contributions from straggler clients [40] or explicitly selecting clients for local training based on their computing capabilities [41, 42, 43]. Nonetheless, these methods are best suited for cross-device FL, where only a subset of clients participates in each training round. They do not align well with cross-silo FL where there are only a few FL clients and ensuring the participation of every client is usually vital for maintaining the robustness of the global model.
@@
Communication heterogeneity is originally rooted in the intermittent availability of client devices due to the limited power and bandwidth in cross-device FL settings, which is less of an issue in cross-silo FL. However, as foundation models increasingly dominate various domains and applications, the interest in using FL to train or fine-tune these models has surged. This surge has led to a substantial increase in communication costs, which become a critical factor affecting the efficiency of FL training [44, 45, 46]. Consequently, improving the efficiency and robustness of transferring large model parameters has become critically important as well[47]. To address this situation, some client selection methods have been proposed to mitigate communication issues in cross-device FL by strategically selecting clients based on their availability, data quality, and performance [48, 49, 50, 51]. Other approaches focus on generic FL settings by applying compression techniques to large model parameters, thereby reducing the overall communication workload [52, 53, 26].
@@ II-B Attacks and Security Concerns in Federated Learning
Gradient inversion algorithms, for example, can reveal information about the private training data by iteratively updating a randomly initialized sample to match its gradient update with the actual model gradient. These algorithms are particularly effective in the early stages of training where the gradients contain more information about the training data [60, 61, 62, 63]. Countermeasures to these inversion attacks include increasing training batch sizes [62], implementing differential privacy techniques to add noise to model gradients [8, 18], compressing model gradients [64, 65], and employing cryptographic strategies such as homomorphic encryption to secure the gradients [66].
Since the FL server cannot access and inspect the client training data, FL is also vulnerable to attacks from Byzantine clients [67], which may either submit corrupted model parameters (model poisoning) or use tampered data for training (data poisoning) [68, 69, 70, 71, 72, 73]. To counter these, several algorithms have been developed to exclude models whose parameters significantly deviate from the norm [74, 75, 70]. Other approaches use the client models to inversely derive training data and then exclude models if their derived data significantly diverges from other models. Alternatively, some solutions assume that the FL server holds a clean and secret validation dataset to evaluate and potentially exclude poorly performing client models from the aggregation process [76, 77].
Beyond algorithmic defenses, addressing malicious attacks in FL can also be achieved through system-level enhancements, particularly by integrating with identity and access management (IAM) services [78]. Such integration enables the creation of secure federations that permit only trusted and known clients to participate in FL experiments, thus fundamentally resolving security concerns at their root.
@@ II-C Existing Federated Learning Frameworks
We conduct a brief survey of twelve popular open-source federated learning frameworks, including the previous version of Appfl without advancements [22] (Appfl-v0), focusing on their approaches to heterogeneity and security challenges, usability, and extensibility for various application scenarios. The results are summarized in Table I.
In addressing data heterogeneity, most frameworks implement advanced client training and server aggregation strategies to mitigate the client drift issue, with the exception of LEAF [54], one of the earliest FL frameworks. Regarding computation heterogeneity, while all frameworks include synchronous FL algorithms, only FedScale [21], FLGo [58], FedLab [59], and Appfl offer asynchronous communication stack and the corresponding asynchronous aggregation strategies. For communication heterogeneity concerns, we evaluate whether the frameworks feature lossless or lossy compression algorithms to reduce the communication loads and whether they provide versatile communication stacks that support multiple protocols, enhancing efficiency and adaptability to different deployment requirements and scenarios.
In terms of the privacy and security challenges, our investigation focuses on whether the frameworks incorporate privacy preservation algorithms such as differential privacy, and integrate IAM services for user authentication and authorization. Most existing frameworks support privacy preservation to some extent, with FLARE [56], OpenFL [57], Flower [19], FedML [20], and Appfl featuring IAM integration for verifying user identities and managing access to specific FL experiments.
As for the usability, most of the frameworks facilitate both the simulation of FL experiments within the same machine or cluster and the real deployment among distributed clients. However, LEAF and TFF [40] are limited to simulation environments only. To evaluate the extensibility and customization capabilities of the frameworks, we access their support for different FL variants beyond the traditional federated learning, specifically vertical federated learning (VFL) [79], hierarchical federated learning (HierFL) [80], and decentralized federated learning (DFL) [81]. FederatedScope [55], FLARE, OpenFL, FLGo, and Flower provide use cases in VFL settings, and FedML and Appfl extend support to all three variants.
III Architecture and Implementation
@@
The Appfl framework is a Python package available on PyPI. Figure 1 provides an overview of its new software architecture. Appfl defines a server agent and a client agent, connected by the communicator, to represent the FL server and clients in performing the primary aggregation task and other necessary tasks for running FL experiments. The server agent is mainly composed of a scheduler module that orchestrates the aggregation of client local models under various synchronicity settings, an aggregator module that aggregates the local models passed from the scheduler to update global model, and a privacy module for additional privacy protection. The client agent consists of a trainer module responsible for training the ML model using the confidential local dataset and a privacy module for the privacy preservation algorithms. The communicator facilitates robust communication between the server and clients, supporting multiple communication protocols for exchanging task control signals and data, with an option to separate the transmission of control signals and data via a data connector. Additionally, the communicator incorporates several compressors for improved efficiency and authenticators for securing the FL experiments. Overall, Appfl incorporates solutions for various challenges in FL and is designed to be modular and extensible, facilitating easy integration of new algorithms and strategies to address FL challenges. The following subsections detail the key components of Appfl.
@@ III-B Communication Stack
For client-driven communication protocols, the Appfl communicator provides a server communicator that defines handlers for various types of requests, such as sending general client configurations and performing global aggregation, by interacting with the server agent. Additionally, a client communicator assists the client agents in sending requests to the server. As for Globus Compute, the server-driven communication protocol, it is a distributed function-as-a-service platform that can dispatch Python functions to run on remote machines. The Appfl communicator provides a Globus Compute server communicator to send various tasks, such as local training, to run on the remote client machines and collect results back for conducting FL experiments. Overall, Appfl supports commonly used server request handlers and client task implementations and provides a user-friendly interface that enables developers to easily define new request handlers or tasks without in-depth knowledge of the underlying communication protocol.
While the communication protocols can transfer the task control signals (i.e., requests in client-driven and tasks in server-driven protocols) along with the associated data, Appfl provides an option to separate the transfer of task controls from the associated model parameters through the integration with ProxyStore [25]. ProxyStore can create a proxy for any target Python object, providing a lightweight reference that can remotely resolve the target object when used. When the proxy is resolved, the object is transferred via an underlying data connector. Appfl currently supports two connectors: an S3 connector, which uses AWS S3 buckets for data transfer, and a ProxyStore endpoint connector, which employs the ProxyStore-hosted relay server. The integration with ProxyStore offers two main benefits: (1) it prevents exceeding the maximum data size limits imposed by certain communication protocols (e.g., Globus Compute restricts task arguments and result sizes to 10 MB to reduce server load, thus making data transfer separation a must when exchanging large model parameters), and (2) it offers users a variety of data transmission options for different communication scenarios and facilitates easy integration of other efficient data transmission methods suitable for their specific use cases to accelerate the FL communication, regardless of the communication protocol in use.
Furthermore, Appfl incorporates a range of data compressors to enhance communication efficiency, crucial for transferring parameters of large models or operating in environments with limited network bandwidth. It supports various lossless compressors including zstd [82], gzip [83], and blosc [84], as well as lossy data compressors including SZ2 [85], SZ3 [86], and ZFP [87]. These compressors can help reduce the communication load, enabling faster data transfer between the server and clients.

### v2

@@ Abstract
Federated learning (FL) is a distributed machine learning paradigm enabling collaborative model training while preserving data privacy. In today’s landscape, where most data is proprietary, confidential, and distributed, FL has become a promising approach to leverage such data effectively, particularly in sensitive domains such as medicine and the electric grid. Heterogeneity and security are the key challenges in FL, however, most existing FL frameworks either fail to address these challenges adequately or lack the flexibility to incorporate new solutions. To this end, we present the recent advances in developing Appfl, an extensible framework and benchmarking suite for federated learning, which offers comprehensive solutions for heterogeneity and security concerns, as well as user-friendly interfaces for integrating new algorithms or adapting to new applications. We demonstrate the capabilities of Appfl through extensive experiments evaluating various aspects of FL, including communication efficiency, privacy preservation, computational performance, and resource utilization. We further highlight the extensibility of Appfl through case studies in vertical, hierarchical, and decentralized FL. Appfl is fully open-sourced on GitHub at https://github.com/APPFL/APPFL.
@@ I Introduction
Availability of extensive training data is becoming increasingly crucial for developing more capable machine learning (ML) models, especially as these models continue to grow in size and complexity. Nonetheless, most of the data in today’s landscape is confidential and distributed across various data silos [1]. This distribution makes it difficult to collect the data for centralized model training, posing significant challenges in fully leveraging the existing data to train more powerful ML models. In this context, federated learning (FL), a distributed ML paradigm, offers a promising solution to utilize data from multiple data owners without direct data sharing [2, 3].
@@
In FL, multiple data owners, referred to as clients, collaborate under a central server to train a shared ML model by iterating two steps: (1) each client trains an ML model using its local dataset and submits the updated model to the server, and (2) the server aggregates these local models to update the global model and then sends it back to the clients for further local training. In this way, FL leverages data from multiple sources to build a more powerful and robust model without data centralization, thereby protecting data privacy. FL has been widely adopted in domains such as medicine [4, 5], finance [6], and electric grid [7], where data privacy is paramount. Depending on the amount, capability, and availability of client devices, FL is broadly categorized into two types, cross-device FL and cross-silo FL [3]. In cross-device FL, numerous mobile or IoT devices with limited computing power and intermittent availability collaboratively train relatively small models such as keyboard suggestion models [8]. In contrast, cross-silo FL involves fewer but more reliable and powerful clients, typically represented by large data silos and institutions, to develop more complex ML models with extensive parameters.
@@
While FL can be conceptually simplified to traditional machine learning with an additional global aggregation operation, its distributed nature introduces significant challenges in terms of heterogeneity and security. Data heterogeneity, stemming from the unbalanced, and non-independent and identically distributed (non-IID) nature of client local datasets, can lead to varied local training objectives across clients and potentially degrade the performance of the global model [9]. Additionally, the heterogeneity in computation and communication, caused by diverse computing capabilities and network connectivity of client devices, can severely impact the efficiency of FL training [10]. This is particularly problematic in synchronous FL algorithms, where the server has to wait for all clients to submit their local models before global aggregation. With regard to security, FL is vulnerable to various attacks. Untrusted clients might maliciously attack FL experiments by submitting corrupted local models, and there is also a risk that training data could be reconstructed from the model updates sent by clients, thereby compromising data privacy [3, 11].
@@
Most existing FL frameworks, such as Flower [12], FedML [13], and FedScale [14], do not adequately address the full spectrum of FL challenges. For example, some frameworks do not support asynchronous aggregation that could improve training efficiency, lack implementations of robust authentication, or fail to offer user-friendly interfaces for easy integration of new algorithms. To bridge these gaps, we developed the Advanced Privacy-Preserving Federated Learning (Appfl) framework, a comprehensive and extensible FL framework that builds on and improves the work presented in [15]. Appfl features advanced aggregation strategies to address data heterogeneity [9] and various asynchronous aggregation strategies to boost training efficiency among heterogeneous computing resources [16]. Additionally, Appfl incorporates versatile communication protocols, data transfer methods, and compression strategies to meet different communication requirements and enhance communication efficiency. It also includes robust authentication via Globus [17], along with plugins for adding new authentication methods, and implements privacy preservation strategies [18] to prevent the reconstruction of training data. Moreover, Appfl is extensible; it follows a modular design with detailed documentation that enables users and developers to seamlessly adapt the framework for different application use cases and integrate custom algorithmic solutions to tackle various FL challenges.
@@
  • Advance Appfl, an open-source and well-documented [^1] The documentation of Appfl is available at https://appfl.ai. FL framework for both FL users and developers that provides established solutions to common FL challenges for FL users and offers flexible and modular interfaces facilitating easy integration of new algorithmic solutions for FL developers
@@ II-A Heterogeneity in Federated Learning
Heterogeneity is one of the key challenges in FL due to its distributed nature. This heterogeneity can be categorized into three primary types: data heterogeneity, computation heterogeneity, and communication heterogeneity.
@@
Data heterogeneity arises from the fact that client datasets are unbalanced and non-IID, meaning they may not be representative of the overall population. This discrepancy leads to varying local training objectives among clients, causing their locally trained models to diverge from one another, a phenomenon known as client drift [19]. As a result, simple weighted averaging of local models, as in the FedAvg strategy [2], may degrade the performance of the global model as data heterogeneity increases [9]. Several solutions have been proposed to address this issue on both the server and client sides. For instance, server-side optimizations such as FedAvgM [9], FedAdam, FedAdagrad, and FedYogi [20] have been introduced to enhance FL performance on non-IID data. Client-side approaches such as SCAFFOLD [19] incorporate correction terms into the client’s local objective function to reduce drift between local and global models. Additionally, selecting participating clients based on data quality and relevance has been explored as well as an effective solution [21, 22, 23].
@@
Computation heterogeneity occurs when the computing devices of FL clients have varying computing power, resulting in large variants in the local training times. This variance poses challenges for synchronous FL strategies, where the server must wait for all clients to submit their local models before global aggregation. Delays from slower clients can reduce the overall training efficiency and lead to underutilization of computing resources. In order to address this issue, various asynchronous aggregation strategies have been proposed. These strategies, including FedAsync [10], FedBuff [24], and FedCompass [16], update the global model immediately upon receiving models from one or a few clients. These methods are beneficial in environments with heterogeneous computing capabilities as they minimize client idle time. Other approaches include disregarding contributions from straggler clients [25] or explicitly selecting clients for local training based on their computing capabilities [26]. Nonetheless, these methods are best suited for cross-device FL, where only a subset of clients participates in each training round. They do not align well with cross-silo FL where there are only a few FL clients and ensuring the participation of every client is vital for maintaining the robustness of the global model.
@@
Communication heterogeneity is originally rooted in the intermittent availability of client devices due to the limited power and bandwidth in cross-device FL, which is less of an issue in cross-silo FL. However, as foundation models increasingly dominate various domains, the interest in using FL to train or fine-tune these models has surged. This surge has led to a substantial increase in communication costs, which become a critical factor affecting the FL training efficiency [27]. Consequently, improving the efficiency and robustness of transferring large model parameters has become critically important as well[28]. To address this situation, some client selection methods have been proposed to mitigate communication issues in cross-device FL by strategically selecting clients based on their availability, data quality, and performance [29]. Other approaches focus on generic FL settings by applying compression or pruning techniques to large model parameters, thereby reducing the communication workload [30, 31].
@@ II-B Attacks and Security Concerns in Federated Learning
Gradient inversion algorithms, for example, can reveal information about the private training data by iteratively updating a randomly initialized sample to match its gradient update with the actual model gradient. These algorithms are particularly effective in the early stages of training where the gradients contain more information about the training data [37]. Countermeasures to these inversion attacks include increasing training batch sizes [38], implementing differential privacy techniques to add noise to model gradients [5, 11], and compressing model gradients [39].
Since the FL server cannot inspect the client training data, FL is also vulnerable to attacks from Byzantine clients, which may either submit corrupted model parameters (model poisoning) or use tampered data for training (data poisoning) [40]. To counter these, several algorithms have been developed to exclude models whose parameters significantly deviate from the norm [41]. Alternatively, some solutions assume that the FL server holds a clean and secret validation dataset to evaluate and potentially exclude poorly performing client models from the aggregation process [42]. Beyond algorithmic defenses, addressing malicious attacks in FL can also be achieved through system-level enhancements, particularly by integrating with identity and access management (IAM) services [43]. Such integration enables the creation of secure federations that permit only trusted and known clients to participate in FL experiments, thus alleviating security concerns at their root.
@@ II-C Existing Federated Learning Frameworks
We conduct a brief survey of several popular open-source federated learning frameworks, including the work we built upon without advancements [15] (denoted as Appfl-v0), focusing on their solutions to heterogeneity and security challenges, usability, and extensibility for various application scenarios. The results are summarized in Table I.
In addressing data heterogeneity, most frameworks implement advanced client training and server aggregation strategies to mitigate client drift issues, with the exception of LEAF [32]. Regarding computation heterogeneity, while all frameworks include synchronous FL algorithms, only FedScale [14], FedLab [36], and Appfl offer asynchronous communication stack and corresponding asynchronous aggregation strategies. For communication heterogeneity concerns, we evaluate whether the frameworks feature lossless or lossy compression algorithms to reduce the communication loads and whether they provide versatile communication stacks that support multiple protocols, enhancing efficiency and adaptability to different deployment requirements and scenarios.
For the privacy and security challenges, our investigation focuses on whether the frameworks incorporate privacy preservation or enhancing algorithms, and integrate IAM services for user authentication and authorization. Most existing frameworks support privacy preservation to some extent, with FLARE [34], OpenFL [35], Flower [12], FedML [13], and Appfl featuring IAM integration for verifying user identities and managing access to specific FL experiments.
As for usability, most of the frameworks facilitate both the simulation of FL experiments within the same machine or cluster and the real deployment among distributed clients. However, LEAF and TFF [25] are limited to simulation environments only. To evaluate the extensibility and ease of customization of the frameworks, we access their support for different FL variants beyond the traditional federated learning, specifically vertical federated learning (VFL) [44], hierarchical federated learning (HierFL) [45], and decentralized federated learning (DFL) [46]. FederatedScope [33], FLARE, OpenFL, and Flower provide use cases in VFL settings, and FedML and Appfl extend support to all three variants.
III Framework Architecture and Implementation
@@
The Appfl framework is a Python package available on PyPI. Figure 1 provides an overview of its new software architecture. Appfl defines a server agent and a client agent, connected by the communicator, to represent the FL server and clients in performing the primary aggregation task and other necessary tasks for running FL experiments. The server agent is mainly composed of a scheduler module that orchestrates the aggregation of client local models under various synchronicity settings, an aggregator module that aggregates the local models passed from the scheduler to update global model, and a privacy module for additional privacy protection. The client agent consists of a trainer module responsible for training the ML model using the confidential local dataset and a privacy module for the privacy preservation algorithms. The communicator facilitates robust communication between the server and clients, supporting multiple communication protocols for exchanging task control signals and data, with an option to separate the transmission of control signals and data via a data connector. Additionally, the communicator incorporates several compressors for improved efficiency and authenticators for securing the FL experiments. Overall, Appfl incorporates solutions for various challenges in FL and is designed to be modular and extensible, facilitating easy integration of new algorithms and strategies to address FL challenges. The following subsections detail several key components of Appfl.
@@ III-B Communication Stack
For client-driven communication protocols, the Appfl communicator provides a server communicator that defines handlers for various types of requests, such as sending configurations and performing global aggregation, by interacting with the server agent. Additionally, a client communicator assists the client agents in sending requests to the server. As for Globus Compute, the server-driven communication protocol, it is a distributed function-as-a-service platform that can dispatch Python functions to run on remote machines. The Appfl communicator provides a Globus Compute server communicator to send various tasks, such as local training, to run on the remote client machines and collect results back for conducting FL experiments. Overall, Appfl supports commonly used server request handlers and client task implementations and provides a user-friendly interface that enables developers to easily define new request handlers or tasks without in-depth knowledge of the underlying communication protocol.
While the communication protocols can transfer the task control signals (i.e., requests in client-driven and tasks in server-driven protocols) along with the associated data, Appfl provides an option to separate the transfer of task controls from the associated model parameters through the integration with ProxyStore [48, 49]. ProxyStore can create a proxy for any target Python object, providing a lightweight reference that can remotely resolve the target object when used. When the proxy is getting resolved, the object is transferred via an underlying data connector from the producer to the consumer. Appfl currently supports two connectors: an S3 connector, which uses AWS S3 buckets for remote data transfer, and a ProxyStore endpoint connector, which transfers data via the ProxyStore-hosted relay server. The integration with ProxyStore offers two main benefits: (1) it prevents exceeding the maximum data size limits imposed by certain communication protocols (e.g., Globus Compute restricts task arguments and result sizes to 10 MB to reduce its service load, thus making data transfer separation a must when exchanging large model parameters), and (2) it offers users a variety of data transmission options for different communication scenarios and facilitates easy integration of other efficient data transmission methods suitable for their specific use cases to accelerate the FL communication, regardless of the communication protocol in use.
Furthermore, Appfl incorporates a range of data compressors to enhance communication efficiency, crucial for transferring parameters of large models or operating in environments with limited network bandwidth. It supports various lossless compressors including zstd, gzip, and blosc, as well as lossy data compressors including SZ2 [50], SZ3 [51], and ZFP [52]. These compressors can help reduce the communication load, enabling faster data transfer between the server and clients.